If peanuts cost $.85 per pound and cashews cost $1.05 per pound, how much of each should be used to make a 12 pound mix that cost $.90?
Answer by htmentor(1343)

If peanuts cost $.85 per pound and cashews cost $1.05 per pound, how much of each should be used to make a 12 pound mix that cost $.90?
================
Let p = the number of pounds of peanuts
Then 12 - p = the number of pounds of cashews
We can write the following expression for the total value of the mixture:
85p + 105(12-p) = 12*90 [in cents]
Solve for p:
85p + 1260 - 105p = 1080
20p = 180
p = 9
So the mix should contain 9 lbs. of peanuts and 3 lbs. of cashews